# Interpolated Stochastic Interventions based on Propensity Scores, Target Policies and Treatment-Specific Costs

### v2

@@ Abstract
We introduce families of stochastic interventions for discrete treatments that connect causal modeling to cost-sensitive decision making. The interventions arise from a cost-penalized information projection of the independent product of the organic propensity and a user-specified target, yielding closed-form Boltzmann–Gibbs couplings. The induced marginals define modified stochastic policies that interpolate smoothly, via a single tilt parameter, from the organic law or from the target distribution toward a product-of-experts limit when all destination costs are strictly positive. One of these families recovers and extends incremental propensity score interventions, retaining identification without global positivity. For inference, we derive efficient influence functions under a nonparametric model for the expected outcomes after these policies and construct one-step estimators with uniform confidence bands. In simulations, the proposed estimators improve stability and robustness to nuisance misspecification relative to plug-in baselines. The framework can operationalize graded scientific hypotheses under realistic constraints: because inputs are modular, analysts can sweep feasible policy spaces, prototype candidates, and align interventions with budgets and logistics before committing experimental resources. This could help close the loop between observational evidence and resource-aware experimental design.
@@ Introduction
Stochastic / soft interventions offer an expressive alternative, capturing hypothetical scenarios in which the treatment assignment mechanism is altered through probabilistic or functional shifts, typically as a function of observed covariates, while the rest of the DGP is left unchanged (Correa and Bareinboim 2020a, b). This formulation enables researchers to define, analyze, and estimate causal effects under interventions that are more adaptive, flexible and aligned with the practical constraints of real-world policy implementation (Haneuse and Rotnitzky 2013; Sarvet et al. 2023).
@@
Mechanistically, IPIs differ from other soft interventions such as general stochastic interventions (GSI) (Correa and Bareinboim 2020a, b) and modified treatment policies (MTP) (Haneuse and Rotnitzky 2013; Díaz et al. 2023), which include shift intervention policies (SIP) (Sani, Lee, and Shpitser 2020). GSIs allow ample changes to the exposure mechanism, for example, reducing the parent set or introducing auxiliary noise, and they may alter the $\sigma$-algebra of its conditional distribution. MTPs set the counterfactual treatment deterministically as a function of the observed treatment and covariates, with SIPs restricting the rule to depend only on the observed treatment. By contrast, IPIs operate solely by tilting the organic propensity score, preserve the underlying measurable structure of the exposure law, and define counterfactual treatments by sampling from the tilted distribution (Kennedy 2018; Bonvini et al. 2023). Recently, generalized policies (GPs) have been proposed, defined as a marginal of an optimal transport map, matching the distribution induced by a reallocation mechanism conditional on covariates, observed treatment, and auxiliary noise. They have been shown to offer favorable properties for partial identification under latent confounding (Levis et al. 2024).
The expected outcome after an IPI can be identified without the global positivity condition required for hard interventions, provided conditional ignorability holds given a backdoor admissible adjustment set. This is important because positivity is often violated in practice, especially with high-dimensional covariates or in longitudinal and dynamic treatment settings (Kennedy 2018; Bonvini et al. 2023). When identifiable, the expected outcome after an IPI can be estimated with efficient semiparametric methods that offer favorable statistical properties, including robustness to misspecification and root-$n$ consistency. Common implementations use either one-step estimators (Bonvini et al. 2023) or targeted minimum loss estimation (TMLE) (Naimi et al. 2021). Inference can be conducted pointwise, at a fixed tilt parameter (for example, a specific odds-ratio shift), or uniformly over a range of values to recover the full intervention response curve (Kennedy 2018).
@@
To address this, we introduce a formulation that yields two families of stochastic interventions governed by a single tilt parameter. The first family smoothly interpolates, in a cost-aware manner, from a non-intervention to the product of experts (PoE) blend of the organic propensity score and a pre-specified target policy; the second interpolates from the target policy to the PoE. The PoE limit arises when all treatment options carry strictly positive costs; if some actions are costless, the limiting behavior depends on the zero-cost set. These cost-aware interpolations are derived from a cost-penalized I-projection and can be viewed as a limiting case of a relaxed optimal transport problem. The first family directly generalizes IPIs, accommodates cost structures and target policies, and, as with IPIs, does not require the global positivity conditions demanded by hard interventions. When the target distribution is non-degenerate, the resulting target marginal also defines a valid stochastic intervention under a positivity condition.
Under standard identification conditions for observational data, we derive efficient influence functions in a nonparametric model for the expected outcome after these interventions, develop robust one-step estimators, and construct asymptotically valid confidence band.
We believe these ideas can help make interventions in causal inference more adaptable and better aligned with practical constraints in domains such as healthcare and economics. They also suggest several directions for future work, including exploring connections with optimal transport.
@@
Let $A\in\mathcal{A}$ be a discrete point-exposure variable, $Y\in\mathbb{R}$ a continuous outcome variable, $W\in\mathcal{W}$ a vector of pre-exposure covariates, and $\pi(a\,|\,w):=\mathbb{P}(A=a\,|\,W=w)$ the propensity score of treatment option $a\in\mathcal{A}$. Potential outcomes $Y^{a}$ encode unit-level counterfactuals after a hard intervention $\operatorname{do}(A=a)$ (Pearl 2009). Under counterfactual consistency, positivity $\pi(a\,|\,w)\in(0,1),\forall w\in\mathcal{W}$, and conditional ignorability / backdoor admissibility of $W$, one can identify the expected outcome after intervention $\operatorname{do}(A=a)$ from observational data via the $g$-computation / backdoor formula, as $\mathbb{E}\left[Y^{a}\right]=\mathbb{E}\left[Y\,|\,\operatorname{do}(A=a)\right]=\mathbb{E}_{W}\left\{Q(Z,a)\right\}$, with $Q(w,a)=\mathbb{E}[Y\,|\,W=w,A=a]$.
@@
For a binary point-exposure $A$, an IPI can be defined as a family of stochastic interventions that replace the organic propensity score $\pi(1\,|\,w)$ by:
@@
IPIs provide a smooth interpolation between a non-intervention and a hard intervention. Specifically, when $\delta=0$, the intervention leaves the propensity score unchanged, i.e., $\widetilde{\pi}_{\delta}(a\,|\,w)={\pi}(a\,|\,w)$. As $\delta\to\infty$, the modified mechanism approaches a deterministic assignment to treatment value $a=1$, i.e., $\widetilde{\pi}_{\delta}(a\,|\,w)\to\mathbb{I}(a=1)$; conversely, as $\delta\to$-$\infty$, the intervention converges to always assigning $a=0$, i.e., $\widetilde{\pi}_{\delta}(a\,|\,w)\to\mathbb{I}(a=0)$.
The expected outcome after a stochastic intervention, such as an IPI with modified exposure mechanism $A\sim\widetilde{\pi}_{\delta}$, can be denoted using notation $\mathbb{E}[Y\fatsemi\sigma_{A}=\widetilde{\pi}_{\delta}]$ (Correa and Bareinboim 2020a, b), by $\mathbb{E}[Y^{\widetilde{\pi}_{\delta}}]$ (Naimi et al. 2021), or by $\mathbb{E}\left[Y\,|\,\operatorname{do}(A\sim\widetilde{\pi}_{\delta})\right]$. Under conditional ignorability / backdoor admissibility of $W$, it is identified as:
@@
Notably, unlike the case of a hard intervention, this causal query remains identifiable without requiring a global positivity condition (Kennedy 2018; Bonvini et al. 2023).
@@
While exponentially tilted distributions have been characterized as standard I-projections under a mean constraint (Levis et al. 2024), and IPIs have been used to model interventions under limited resource and treatment costs (Sarvet et al. 2023), a broader connection between IPIs and limiting cases of relaxed optimal transport problems that accommodate cost structures and target distributions has not been described. In this work, we provide that characterization.
@@
Let $A$ be a categorical point-exposure variable with domain $\mathcal{A}=\{\alpha_{1},\dots,\alpha_{K}\}$, where the $K$ treatment options may potentially include a placebo or null-treatment. Let $\nu$ be a target marginal distribution over $\mathcal{A}$ and let $\pi(a\,|\,w)$ be the propensity score of treatment $A=a$ given pre-exposure covariate profile $W=w$. Let $c(a^{\prime},a^{\prime\prime})$ be the cost of reallocating a unit from treatment $A=a^{\prime}$ to treatment $A=a^{\prime\prime}$, which does not depend on the profile $w$. Then, for all $a\in\mathcal{A}$ and $w\in\mathcal{W}$,
$\pi^{\star}_{\delta}(a\,|\,w)$ $:=\frac{\pi(a\,|\,w)\sum_{a^{\prime\prime}\in\mathcal{A}}\nu(a^{\prime\prime})\,e^{-\delta c(a,a^{\prime\prime})}}{\sum_{a^{\prime},a^{\prime\prime}\in\mathcal{A}}\pi(a^{\prime}\,|\,w)\,\nu(a^{\prime\prime})\,e^{-\delta c(a^{\prime},a^{\prime\prime})}},$ (5)
$\nu^{\star}_{\delta}(a\,|\,w)$ $:=\frac{\nu(a)\sum_{a^{\prime}\in\mathcal{A}}\pi(a^{\prime}\,|\,w)\,e^{-\delta c(a^{\prime},a)}}{\sum_{a^{\prime},a^{\prime\prime}\in\mathcal{A}}\pi(a^{\prime}\,|\,w)\,\nu(a^{\prime\prime})\,e^{-\delta c(a^{\prime},a^{\prime\prime})}}$ (6)
are the marginal distributions of the CPIP solution in (4) when the source is $\pi(\cdot\,|\,w)$ and the target is $\nu$. We refer to these as the tilted source / target marginal distributions. Derivation is provided in the technical appendix A1.
@@
    the target marginal $\nu$ be the degenerate distribution that always assigns treatment, $\nu(a)=\mathbb{I}(a=1)$,
@@
The CPIP objective is strictly convex and well posed for $\delta>0$, yielding a unique solution. For $\delta<0$, the joint distribution $\gamma^{\star}_{\delta}$ in (4) is not the optimizer of the CPIP program; however, when the action set $\mathcal{A}$ is finite and costs are bounded, the associated Boltzmann–Gibbs kernel is summable, so $\gamma^{\star}_{\delta}$ remains normalizable and produces smooth and closed-form parametric extensions for $\pi^{\star}_{\delta}$ and $\nu^{\star}_{\delta}$. In this extrapolative regime, the induced tilted marginals shift mass toward higher-cost and repulsive pairings and are pushed away from the independence law $\pi(\cdot\,|\,w)\otimes\nu$, favoring more adversarial couplings. For continuous spaces with potentially unbounded costs, additional integrability or tail conditions are required for $\delta<0$ to deliver well posed stochastic policies.
Notably, within the setting of remark 1, the tilted target marginal coincides with the input target distribution, so $\nu^{\star}_{\delta}(a\,|\,w)=\nu(a)=\mathbb{I}(a=1)$ for all $\delta>0$ and all $w\in\mathcal{W}$. Consequently, it does not provide a basis for a differentiated stochastic policy. In the next remark, we introduce a generalization of IPIs based on stochastic plans that incorporates treatment-specific costs and a flexible target policy.
@@
When all treatment costs are positive, both reduce to the product of experts (PoE) distribution $\operatorname{PoE}(a)\propto\pi(a\,|\,w)\,\nu(a)$ (Hinton 1999).
Although, as noted in remark 1, typical incremental IPIs are defined using the tilted source distribution $\pi^{\star}_{\delta}$, we argue that the tilted target distribution $\nu^{\star}_{\delta}$ can also be used to define stochastic policies, particularly when the target $\nu$ is non-degenerate. For example, suppose the ideal policy prescribes treatment 1 to 80% of units, treatment 2 to 10%, and no treatment to the remaining 10%. Then, in the case with positive costs, $\nu^{\star}_{\delta}$ provides a smooth interpolation between this ideal allocation ($\delta=0$) and the PoE law ($\delta\to\infty$).
@@
Since the CPIP objective imposes no soft or hard constraints on the marginals, the optimizer $\gamma^{\star}_{\delta}$ in (4) is not a relaxed optimal transport plan from $\pi$ to $\nu$. Consequently, although $\nu^{\star}_{\delta}$ provides a smooth, cost-aware, one-parameter deformation of the input distributions, it should not be interpreted as the pushforward of $\pi$ through $\gamma^{\star}_{\delta}$. One may instead enforce the source marginal to be equal or close to $\pi$ while leaving the other marginal unconstrained by adding marginal penalties, as in relaxed optimal transport (Frogner et al. 2015; Chizat et al. 2018a, b), but such formulations typically do not admit closed-form solutions. As an operational alternative, one can consider the pushforward distribution, obtained by applying the Markov kernel induced by $\gamma^{\star}_{\delta}$ to the source law:
$\nu^{\star\star}_{\delta}(a\,|\,w)\;:=\;\sum_{a^{\prime}\in\mathcal{A}}\frac{\gamma^{\star}_{\delta}(a^{\prime},a\,|\,w)}{\pi^{\star}_{\delta}(a^{\prime}\,|\,w)}\,\pi(a^{\prime}\,|\,w),$ (13)
which is well defined because $\pi^{\star}_{\delta}(a\,|\,w)$ is proportional to $\pi(a\,|\,w)$ with a strictly positive weight (hence $\pi(a\,|\,w)/\pi^{\star}_{\delta}(a\,|\,w)>0$ for all $a$ and $w$). Here $\nu^{\star\star}_{\delta}(\cdot\,|\,w)$ represents the distribution induced by applying the stochastic plan encoded in $\gamma^{\star}_{\delta}$ to units drawn from the organic source $\pi(\cdot\,|\,w)$. A systematic comparison between marginal-unconstrained, marginal-penalized, and pushforward formulations is left for future work.
@@
Recent work has extended semiparametric and robust estimation to queries under stochastic interventions (Díaz and van der Laan 2012; Duong, Li, and Xu 2021; Bonvini et al. 2023; Díaz et al. 2023). Semiparametric efficient approaches are prized for accommodating data-adaptive learning while attaining optimal asymptotics under flexible DGP assumptions (Díaz 2019). Many estimators enjoy double or multiple robustness, remaining consistent despite some nuisance misspecification (Daniel 2018). Core frameworks include one-step (Newton–Raphson) corrections (Pfanzagl 1982; Bickel et al. 1993; Robins, Rotnitzky, and Zhao 1995; Robins and Rotnitzky 1997), targeted minimum loss estimation (TMLE) (van der Laan and Rose 2011, 2018), and debiased machine learning (DML) (Chernozhukov et al. 2018). These methods leverage the efficient influence function (EIF) of the smooth functional defining the parameter (Hines et al. 2022).
@@
Let $\mathcal{S}_{\delta}[P]$ denote the functional that evaluates expression (14) at an arbitrary distribution $P$. Similarly, let $\mathcal{T}_{\delta}[P]$ represent the corresponding $P$-functional from expression (15). Let $O_{i}=(W_{i},A_{i},Y_{i})$ denote a sample drawn from the true DGP $P^{*}$. Under a nonparametric model, suitable smoothness and regularity conditions, the uncentered EIF of $\mathcal{S}_{\delta}[P]$ at $P^{*}$, evaluated at point $O_{i}$, exists and is:
@@
Analogously, under a nonparametric model and typical smoothness and regularity conditions, the uncentered EIF of $\mathcal{T}_{\delta}[P]$ at $P^{*}$, evaluated at point $O_{i}$, exists and is given by:
@@
Derivation is provided in technical appendix A4.
For binary exposure $A$ and target distribution $\nu(a)=\mathbb{I}(a=1)$, it is straightforward to verify that $D_{\delta}^{{S}}(O_{i})$ coincides with the uncentered EIF for the expected outcome after an IPI, as derived in prior work (Kennedy 2018; Bonvini et al. 2023).
Note that $\varrho_{0}(a)=0$ and $\pi^{\star}_{0}(a\,|\,w)=\pi(a\,|\,w)$ for all $a\in\mathcal{A},w\in\mathcal{W}$. Hence, when $\delta=0$ and $\nu(a)=\mathbb{I}(a=a^{\prime})$, we recover the standard uncentered EIF for the expected outcome after the hard intervention $\operatorname{do}(A=a^{\prime})$, identified by $\mathbb{E}_{Z}\{Q(Z,a^{\prime})\}$.
Consequently, one-step estimators of the expected outcomes in (14) and (15) are:
@@
using samples $\{O_{i}\}_{i=1}^{n}\overset{iid}{\sim}P^{*}$, and where $\widehat{D}^{S}_{\delta}(\cdot)$ and $\widehat{D}^{T}_{\delta}(\cdot)$ are obtained by plugging the estimated nuisance components $\widehat{\pi}$ and $\widehat{Q}$ into ${D}^{S}_{\delta}(\cdot)$ and ${D}^{T}_{\delta}(\cdot)$, respectively. These estimators eliminate the first-order bias in the von Mises expansion of the corresponding functionals around the true distribution (Pfanzagl 1982; Bickel et al. 1993), even under partial misspecification of some components of the DGP.

### v3

@@ Abstract
We introduce two families of stochastic interventions with discrete treatments that connect causal modeling to cost-sensitive decision making. The interventions arise from a cost-penalized information projection of the independent product of the organic propensity scores and a reference policy, yielding closed-form Boltzmann–Gibbs couplings. The induced marginals define modified stochastic policies that interpolate smoothly, via a tilt parameter, from the organic law or from the reference law toward a product-of-experts limit when all destination costs are strictly positive. The first family recovers and extends incremental propensity score interventions, retaining identification without global positivity. For inference on the expected outcomes after these policies, we derive the efficient influence functions under a nonparametric model and construct one-step estimators. In simulations, the proposed estimators improve stability and robustness to nuisance misspecification relative to plug-in baselines. The framework can operationalize graded scientific hypotheses under realistic constraints. Because inputs are modular, analysts can sweep feasible policy spaces, prototype candidates, and align interventions with budgets and logistics before committing experimental resources.
Extended versionhttp://arxiv.org/abs/2511.11353
@@ Introduction
Stochastic / soft interventions offer an expressive alternative, capturing hypothetical scenarios in which the treatment assignment mechanism is altered through probabilistic or functional shifts, typically as a function of observed covariates, while the rest of the DGP is left unchanged (Correa and Bareinboim 2020a, b). This formulation enables researchers to define, analyze, and estimate causal effects under interventions that are more adaptive, flexible, and aligned with the practical constraints of real-world policy implementation (Haneuse and Rotnitzky 2013; Sarvet et al. 2023).
@@
Mechanistically, IPIs differ from other soft interventions such as general stochastic interventions (GSI) (Correa and Bareinboim 2020a, b) and modified treatment policies (MTP) (Haneuse and Rotnitzky 2013; Díaz et al. 2023), which include shift intervention policies (SIP) (Sani, Lee, and Shpitser 2020). GSIs allow ample changes to the exposure mechanism, for example, reducing the parent set or introducing auxiliary noise, and they may alter the $\sigma$-algebra of its conditional distribution. MTPs set the counterfactual treatment deterministically as a function of the observed treatment and covariates, with SIPs restricting the rule to depend only on the observed treatment. By contrast, IPIs operate solely by tilting the organic propensity score and preserve the underlying measurable structure of the exposure law (Kennedy 2018; Bonvini et al. 2023). Recently, generalized policies (GPs) have been proposed, defined as a marginal of an optimal transport map, matching the distribution induced by a reallocation mechanism conditional on covariates, observed treatment, and auxiliary noise. They have been shown to offer favorable properties for partial identification under latent confounding (Levis et al. 2024).
The expected outcome after an IPI can be identified without the global positivity condition required for hard interventions, provided conditional ignorability holds with a given a backdoor admissible adjustment set. This is important because positivity is often violated in practice, especially with high-dimensional covariates or in longitudinal and dynamic treatment settings (Kennedy 2018; Bonvini et al. 2023). When identifiable, the expected outcome after an IPI can be estimated with efficient semiparametric methods that offer favorable statistical properties, including robustness to misspecification and root-$n$ consistency. Common implementations use either one-step estimators (Bonvini et al. 2023) or targeted minimum loss estimation (TMLE) (Naimi et al. 2021). Inference can be conducted pointwise, at a fixed tilt parameter (for example, a specific odds-ratio shift), or uniformly over a range of values to recover the full intervention response curve (Kennedy 2018).
@@
To address this, we introduce a formulation that yields two families of stochastic interventions indexed by a single tilt parameter. The first family interpolates smoothly, in a cost-aware manner, from the organic law toward its product-of-experts (PoE) blend with a pre-specified reference / target policy, while the second interpolates from the reference policy toward the PoE. The PoE limit is valid when all treatment levels carry strictly positive costs; if some actions are costless, the limiting behavior is determined by the associated zero-cost subset. Both interpolations arise as the marginals of the bivariate distribution that solves a cost-penalized information projection. The first family directly generalizes IPIs to incorporate explicit cost structures and target policies and, as in the IPI framework, avoids the global positivity condition for identification. In addition, when the reference distribution is non-degenerate, the induced target marginal itself defines a valid stochastic intervention under an appropriate positivity condition.
For estimation and inference of mean outcomes after these interpolated policies, we derive the efficient influence functions in a nonparametric model and develop one-step estimators that are robust to outcome model misspecification. We further construct asymptotically valid confidence intervals and uniform confidence bands over a specified segment of the interpolation range.
These ideas support the evaluation of hypotheses that depart gradually from the status quo while explicitly accounting for costs, enabling analysts to prototype candidate policies and align interventions with budgetary and logistical constraints before committing experimental resources, with relevance in domains such as healthcare and economics.
@@
Let $A\in\mathcal{A}$ be a discrete point-exposure, $Y\in\mathbb{R}$ a continuous outcome, $W\in\mathcal{W}$ a vector of pre-exposure covariates, and $\pi(a\,|\,w):=\mathbb{P}(A=a\,|\,W=w)$ the propensity score of treatment option $a\in\mathcal{A}$. Potential outcomes $Y^{a}$ encode unit-level counterfactuals after a hard intervention $\operatorname{do}(A=a)$ (Pearl 2009). Under counterfactual consistency, positivity $\pi(a\,|\,w)\in(0,1),\forall w\in\mathcal{W}$, and conditional ignorability / backdoor admissibility of $W$, one can identify the expected outcome after intervention $\operatorname{do}(A=a)$ from observational data via the $g$-computation / backdoor formula, as $\mathbb{E}\left[Y^{a}\right]=\mathbb{E}\left[Y\,|\,\operatorname{do}(A=a)\right]=\mathbb{E}_{W}\left\{Q(W,a)\right\}$, with $Q(w,a)=\mathbb{E}[Y\,|\,W=w,A=a]$.
@@
For a binary point-exposure $A$, an IPI can be defined as a parametrized family of stochastic interventions that replace the organic propensity score $\pi(1\,|\,w)$ by:
@@
IPIs provide smooth interpolations between a non-intervention and a hard intervention. Specifically, when $\delta=0$, the intervention leaves the propensity score unchanged, i.e., $\widetilde{\pi}_{\delta}(a\,|\,w)={\pi}(a\,|\,w)$. As $\delta\to\infty$, the modified mechanism approaches a deterministic assignment to treatment value $a=1$, i.e., $\widetilde{\pi}_{\delta}(a\,|\,w)\to\mathbb{I}(a=1)$; conversely, as $\delta\to$-$\infty$, the intervention converges to always assigning $a=0$, i.e., $\widetilde{\pi}_{\delta}(a\,|\,w)\to\mathbb{I}(a=0)$.
The expected outcome after a stochastic intervention, such as an IPI with modified exposure mechanism $A\sim\widetilde{\pi}_{\delta}$, can be denoted using notation $\mathbb{E}[Y^{\widetilde{\pi}_{\delta}}]$. Under conditional ignorability / backdoor admissibility of $W$, it is identified by:
@@
Notably, unlike the case of a hard intervention, this causal query does not require a global positivity condition (Kennedy 2018; Bonvini et al. 2023).
@@
While exponentially tilted distributions have been characterized as standard I-projections under a mean constraint (Levis et al. 2024), and IPIs have been used to model interventions under limited resource and treatment costs (Sarvet et al. 2023), a broader connection between IPIs and bivariate I-projections that accommodate cost structures and target distributions has not been described. In this work, we provide that characterization.
@@
Let $A$ be a categorical point-exposure variable with domain $\mathcal{A}=\{\alpha_{1},\dots,\alpha_{K}\}$, where the $K$ treatment options may potentially include a placebo or null-treatment. Let $\nu(\cdot\,|\,w)$ be a target distribution over $\mathcal{A}$ and let $\pi(\cdot\,|\,w)$ be the propensity scores given pre-exposure covariate profile $W=w$. Let $c(a^{\prime},a^{\prime\prime})$ be the cost of reallocating a unit from treatment $A=a^{\prime}$ to treatment $A=a^{\prime\prime}$, which does not depend on the profile $w$. Then, for all $a\in\mathcal{A}$ and $w\in\mathcal{W}$,
$\pi^{\star}_{\delta}(a\,|\,w)$ $:=\frac{\pi(a\,|\,w)\sum_{a^{\prime\prime}\in\mathcal{A}}\nu(a^{\prime\prime}\,|\,w)\,e^{-\delta c(a,a^{\prime\prime})}}{\sum_{a^{\prime},a^{\prime\prime}\in\mathcal{A}}\pi(a^{\prime}\,|\,w)\,\nu(a^{\prime\prime}\,|\,w)\,e^{-\delta c(a^{\prime},a^{\prime\prime})}},$ (5)
$\nu^{\star}_{\delta}(a\,|\,w)$ $:=\frac{\nu(a\,|\,w)\sum_{a^{\prime}\in\mathcal{A}}\pi(a^{\prime}\,|\,w)\,e^{-\delta c(a^{\prime},a)}}{\sum_{a^{\prime},a^{\prime\prime}\in\mathcal{A}}\pi(a^{\prime}\,|\,w)\,\nu(a^{\prime\prime}\,|\,w)\,e^{-\delta c(a^{\prime},a^{\prime\prime})}}$ (6)
are the marginal distributions of the CPIP solution in (4) with source $\pi$ and target $\nu$. We refer to these as the tilted source / target marginal distributions. Derivations are provided in the technical appendix A1.
@@
    the target marginal $\nu$ be the degenerate distribution that always assigns treatment, $\nu(a\,|\,w)=\mathbb{I}(a=1)$,
@@
The CPIP objective is strictly convex and well posed for $\delta\geq 0$, yielding a unique solution. For $\delta<0$, the joint distribution $\gamma^{\star}_{\delta}$ in (4) is not the optimizer of the CPIP program; however, when the action set $\mathcal{A}$ is finite and costs are bounded, the associated Boltzmann–Gibbs kernel is summable, so $\gamma^{\star}_{\delta}$ remains normalizable and produces smooth and closed-form parametric extensions for $\pi^{\star}_{\delta}$ and $\nu^{\star}_{\delta}$. In this extrapolative regime, the induced tilted marginals shift mass toward higher-cost and repulsive pairings and are pushed away from the independence law $\pi\otimes\nu$, favoring more adversarial couplings. For continuous spaces with potentially unbounded costs, additional integrability or tail conditions are required for $\delta<0$ to deliver well posed stochastic policies.
Notably, within the setting of remark 1, the tilted target marginal coincides with the input target distribution, so $\nu^{\star}_{\delta}(a\,|\,w)=\mathbb{I}(a=1)$ for all $\delta\geq 0$ and all $w\in\mathcal{W}$. Consequently, it does not yield a differentiated stochastic policy. In the next remark, we introduce a generalization of IPIs that accommodates treatment-specific costs and a flexible target policy. Although we state the results for an unconditional reference / target $\nu(a)$, the extension to a conditional target $\nu(a\,|\,w)$ is straightforward.
@@
When all treatment costs are positive, both reduce to the product of experts (PoE) distribution $\operatorname{PoE}(a)\propto\pi(a\,|\,w)\,\nu(a)$ (Hinton 1999). This PoE law can be interpreted as a consensus distributional blend of inputs $\pi,\nu$.
We argue that the tilted target distribution $\nu^{\star}_{\delta}$ can also be used to define stochastic policies, particularly when the target $\nu$ is non-degenerate. For example, suppose the ideal policy prescribes treatment 1 to 80% of units, treatment 2 to 10%, and no treatment to the remaining 10%. Then, in the case with positive costs, the tilted target marginal $\nu^{\star}_{\delta}$ provides a smooth interpolation between this ideal allocation ($\delta=0$) and the PoE law ($\delta\to\infty$).
@@
The CPIP criterion optimizes over joint laws $\gamma(\,\cdot,\cdot\,|\,w)$ without imposing or penalizing any marginal constraint. Therefore the solution $\gamma^{\star}_{\delta}$ in (4) should not be viewed as a transport plan that maps $\pi$ to $\nu$. In particular, the tilted marginal $\nu^{\star}_{\delta}(\cdot\,|\,w)$ is simply the second marginal of $\gamma^{\star}_{\delta}$ and does not represent the distribution obtained by pushing $\pi(\cdot\,|\,w)$ through $\gamma^{\star}_{\delta}$. If one wants a transport-style interpretation, one can modify the objective by enforcing the marginals to equal or remain close to the inputs via penalties, as in relaxed optimal transport (Frogner et al. 2015; Chizat et al. 2018a, b). Such formulations, however, typically sacrifice the closed-form solution. As an operational alternative within the CPIP construction, we can define a pushforward of the original source law by using the Markov kernel induced by $\gamma^{\star}_{\delta}$:
$\nu^{\sharp}_{\delta}(a\,|\,w)\;:=\;\sum_{a^{\prime}\in\mathcal{A}}\frac{\gamma^{\star}_{\delta}(a^{\prime},a\,|\,w)}{\pi^{\star}_{\delta}(a^{\prime}\,|\,w)}\,\pi(a^{\prime}\,|\,w),$ (13)
This is well defined provided $\pi^{\star}_{\delta}(a^{\prime}\,|\,w)>0$ when $\pi(a^{\prime}\,|\,w)>0$. For profile-independent destination costs $c(a^{\prime},a^{\prime\prime}\,|\,w)=c(a^{\prime\prime})\,\mathbb{I}(a^{\prime}\neq a^{\prime\prime})$, the kernel does not depend on $a^{\prime}$, and the pushforward reduces to an exponential tilting of $\nu$, i.e., $\nu^{\sharp}_{\delta}(a\,|\,w)\propto\nu(a)\,e^{-\delta c(a)}$. A systematic comparison between the CPIP marginal, marginal-penalized OT variants, and the pushforward is left for future work.
@@
Recent work has leveraged semiparametric efficient methods to estimate causal queries after stochastic interventions (Díaz and van der Laan 2012; Duong, Li, and Xu 2021; Bonvini et al. 2023; Díaz et al. 2023). Semiparametric efficient approaches are prized for accommodating data-adaptive learning while attaining optimal asymptotics under flexible DGP assumptions (Díaz 2019). Many estimators enjoy double or multiple robustness, remaining consistent despite some nuisance misspecification (Daniel 2018). Core frameworks include Newton–Raphson one-step corrections (Pfanzagl 1982; Bickel et al. 1993; Robins, Rotnitzky, and Zhao 1995; Robins and Rotnitzky 1997), targeted minimum loss estimation (TMLE) (van der Laan and Rose 2011, 2018), and debiased machine learning (DML) (Chernozhukov et al. 2018). These methods leverage the efficient influence function (EIF) of the smooth functional defining the parameter (Hines et al. 2022).
@@
Let $\mathcal{S}_{\delta}[P]$ denote the functional that evaluates expression (14) at an arbitrary distribution $P$. Similarly, let $\mathcal{T}_{\delta}[P]$ represent the corresponding $P$-functional from expression (15). Let $O_{i}=(W_{i},A_{i},Y_{i})$ denote a sample drawn from the true DGP $P^{\circ}$. Under a nonparametric model, suitable smoothness and regularity conditions, the uncentered EIF of $\mathcal{S}_{\delta}[P]$ at $P^{\circ}$, evaluated at point $O_{i}$, exists and is:
@@
Analogously, under a nonparametric model and typical smoothness and regularity conditions, the uncentered EIF of $\mathcal{T}_{\delta}[P]$ at $P^{\circ}$, evaluated at point $O_{i}$, exists and is given by:
@@
For binary exposure $A$ and target distribution $\nu(a)=\mathbb{I}(a=1)$, it is straightforward to verify that $D_{\delta}^{{S}}(O_{i})$ coincides with the uncentered EIF for the expected outcome after an IPI, as derived in prior work (Kennedy 2018).
Note that $\varrho_{0}(a)=0$ and $\pi^{\star}_{0}(a\,|\,w)=\pi(a\,|\,w)$ for all $a\in\mathcal{A},w\in\mathcal{W}$. Hence, when $\delta=0$ and $\nu(a)=\mathbb{I}(a=a^{\prime})$, we recover the standard uncentered EIF for the expected outcome after the hard intervention $\operatorname{do}(A=a^{\prime})$.
Consequently, Newton–Raphson one-step estimators of the expected outcomes in (14) and (15) are given by:
@@
using sample $\{O_{i}\}_{i=1}^{n}\overset{iid}{\sim}P^{\circ}$. Here $\widehat{D}^{S}_{\delta}(\cdot)$ and $\widehat{D}^{T}_{\delta}(\cdot)$ are formed by substituting the nuisance estimates $\widehat{\pi}$ and $\widehat{Q}$ into (17) and (18), respectively. These estimators remove the first-order bias in the von Mises expansion of the associated functionals around the true distribution (Pfanzagl 1982; Bickel et al. 1993), are robust to outcome-model misspecification through $\widehat{Q}$, and are asymptotically linear under standard regularity conditions provided the propensity score estimator satisfies a convergence rate of $\norm{\widehat{\pi}-\pi}_{L^{2}(P)}=o_{P^{\circ}}(n^{-1/4})$, or faster.
Derivation and proof are provided in appendix A4 and A5.